zarkam21 (zarkam21):

Two perpendicular lines must: A. have slopes that are negative reciprocals B. cross each other C. cross each other with the same slope D.have slopes that are reciprocals

OpenStudy (aaronandyson):

Is this related to slopes?

zarkam21 (zarkam21):

Yes

OpenStudy (aaronandyson):

Let the two slopes be m_1 and m_2

OpenStudy (aaronandyson):

When these two lines are parallel m_1 = m_2

OpenStudy (jadeishere):

Aaron, it's not discussing parallel lines, actually, quite the oppposite

OpenStudy (aaronandyson):

I need to explain this to further explain it ..in my way XD

OpenStudy (precal):

A

zarkam21 (zarkam21):

So A ?

OpenStudy (precal):

one is positive and one is negative

zarkam21 (zarkam21):

Can everyone confirm this? :)

OpenStudy (aaronandyson):

So when these two lines are perpendicular , m1*m2 = -1

OpenStudy (precal):

reciprocals as well

zarkam21 (zarkam21):

@Jadeishere is A correct?

OpenStudy (precal):

m=1/3 and m=-3 will be perpendicular

OpenStudy (precal):

you could graph it to see that this is true

OpenStudy (aaronandyson):

A looks good.

OpenStudy (kinged):

It actually looks stunning, for some reason...

OpenStudy (jadeishere):

Yes, Zarkam, A is correct. When a line has a negative reciprocal slope, they intersect at exactly a 90 degree angle, basically, flipping the line over the x or the y axis